OpenStudy (anonymous):

Could someone help me with exponential/logarithmic functions? I have a rough idea of what I'm doing, I just need some guidance.

OpenStudy (anonymous):

I'm supposed to answer various questions about buying a car using the formula:\[A(t) = P(1+\frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

is this the same problem as before?

OpenStudy (anonymous):

Forgive me if I sound rude, but may I ask what you mean?

OpenStudy (anonymous):

is this about a car loan for a chevy volt?

OpenStudy (anonymous):

No, this is just for an ordinary car with a value of $15,000 and an interest rate of 6%...

OpenStudy (anonymous):

ok what is the question? btw this is not the formula for any loan this is the formula for compound interest

OpenStudy (anonymous):

for example if you invest \(\%1000\) and \(5\%\) interest compounded monthly, then after \(t\) years you will have \[P(t)=1000(1+\frac{.05}{12})^{12t}\] dollars

OpenStudy (anonymous):

oops i mean if you invest \(\$1000\) !

OpenStudy (anonymous):

That's alright! I do have a quick question. I notice you used P(t) instead of A(t). Do they need to be flipped around? I do apologize for asking so many questions. I get confused somewhat easily when it comes to Math of any kind.

OpenStudy (anonymous):

no it makes no difference you could say \(P(t)\) or \(A(t)\) or \(\Xi(t)\) whatever

OpenStudy (anonymous):

what is the question for this formula?

OpenStudy (anonymous):

I need to find the function in order to know the compounded monthly car payment.

OpenStudy (anonymous):

What I've got so far is: \[A(t) = 15,000(1+ 0.005)^{12t}\]

OpenStudy (anonymous):

ok the formula for monthly payments are not at all like this this is the formula for compound interest weird that this is the second problem in the last half hour that have this mistake something weird is going on

OpenStudy (anonymous):

Does that look right to you, or did I do something wrong?

OpenStudy (anonymous):

what you did was correct for the formula you are using, but i have no idea how you are going to find monthly payments from it

OpenStudy (anonymous):

I don't know either, and that's what's confusing me.

OpenStudy (anonymous):

\[A(t) = 15,000(1+ 0.005)^{12t}\] is the same as \[A(t) = 15,000(1.005)^{12t}\]

OpenStudy (anonymous):

you can't

OpenStudy (anonymous):

I don't need anyone to tell me the answer directly because I want to find it myself, but I don't know how to find the answer.

OpenStudy (anonymous):

the payment formula is much more complicated and completely different

OpenStudy (anonymous):

May I ask how much different?

OpenStudy (anonymous):

the formula you are using \[A(t)=A_0(1+\frac{r}{n})^{nt}\] is for compound interest

OpenStudy (anonymous):

the last formula on the page i just sent is the formula for payments

OpenStudy (anonymous):

Number 2 is roughly the one I'm working with now, but I haven't seen the first one at all.

OpenStudy (anonymous):

the first one is the same as the second if \(n=1\) i.e. once a year

OpenStudy (anonymous):

if \(n=1\) then \[A_0(1+\frac{r}{n})^{nt}=A_0(1+r)^t\]

OpenStudy (anonymous):

Thank you so much! I appreciate your help, and your patience. :) Medal for you, my friend.

OpenStudy (anonymous):

thanks, but i am afraid i have not helped at all still not sure what the question is

OpenStudy (anonymous):

It's alright. :) I'm really confused about it myself, but you tried to help me, and that's something I'm very grateful for. When I figure it out and find out if it's correct or not, I'll let you know. Thanks again!